Circular No. 8237

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

1990 OS S. J. Ostro, Jet Propulsion Laboratory; M. C. Nolan, National As-

tronomy and Ionosphere Center; L. A. M. Benner and J. D. Giorgini, Jet Propulsion Laboratory; J. L. Margot, University of California, Los Angeles; and C. Magri, University of Maine, Farmington, report that Arecibo radar delay-Doppler images (2380 MHz, 12.6 cm) obtained on Nov. 5 and 6 show that 1990 OS (cf. IAUC 5064; MPC 49636) is a binary system. Preliminary estimates of average diameters, based on range extents at 15-m resolution, are 300 and 45 m. The components’ delay-Doppler dispersions suggest spin periods of several hours. The maximum range separation observed is more than 600 m. The pattern of motion suggests an orbital period between 18 and 24 hr. Optical lightcurve observations of this object would be valuable.

SUPERNOVAE 2003ji–2003ju R. Covarrubias, University of Washington, on behalf of the ESSENCE

project (cf. IAUC 8021, 8034), reports the discovery of additional supernovae on multiple R- and I-band images taken at the Blanco 4-m telescope during Oct. 18–31 by project members (plus A. Alda and J. L. Prieto). Spectroscopy was obtained on Oct. 22–31 by P. Berlind, S. Blondin, B. Leibundgut, and T. Matheson with the Very Large Telescope, Gemini telescope, and Mt. Hopkins 1.52-m reflector. The table includes the measured redshift and the estimated age (in days) at the specified discovery date.

SN  2003 UT  α_{2000}  δ_{2000}  R  z Type Age
2003ji Oct. 20 2^{h}07^{m}54.^{s}84 − 3^{o}28′28′′.4 20.9 0.2 Ia  + 15
2003jj Oct. 22 1 07 58.52 + 0 03 01.9 23.1 0.6 Ia  − 1
2003jk Oct. 22 1 08 06.17 + 0 02 20.6 22.5 0.1 Ia  − 11
2003jl Oct. 22 2 28 28.56 − 8 08 44.7 22.4 0.4 Ia  + 7
2003jm Oct. 22 2 28 50.93 − 9 09 58.1 22.9 0.5 Ia  0
2003jn Oct. 22 2 29 21.21 + 9 02 15.6 20.8 0.3 Ia  − 3
2003jo Oct. 22 23 25 24.03 − 9 26 00.6 22.8 0.5 Ia  + 3
2003jp Oct. 22 23 26 03.28 − 8 59 22.7 22.9 0.3 Ib/c +100
2003jq Oct. 22 23 30 51.19 − 9 28 34.0 19.8 0.2 Ia  + 21
2003jr Oct. 24 1 11 06.23 + 0 13 44.2 21.9 0.3 Ia  + 9
2003js Oct. 24 2 29 52.15 − 8 32 28.1 22.0 0.4 Ia  − 5
2003jt Oct. 24 2 31 54.60 − 8 35 48.4 22.0 0.4 Ia  + 7
2003ju Oct. 24 23 27 01.71 − 9 24 04.5 21.6 0.2 Ia  0
2003 November 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green